On the basis of stoicheometry, justify the answer why 1 mL of 0.025 M of Na2S2O3 is equivalent to 1.0 mg/L of DO? Calculate the DO value if your 250 mL sample consumes 18 mL of 0.014 M Na2S2O3?
Answer #1

1.0 ml of 0.025 N thiosulfate is equivalent to 25*10-6 equivalents of iodine which is in turn equivalent to 25*10-6 equivalents of oxygen
25*10-6 equivalents of O2 = 200 * 10-6 g O2

Calculation:

Volume of sodium thiosulfate V1=18 ml

Normality of Sodium thiosulfate N1=0.014M

Volume of Sample V2 = 250 ml

Normality of Titrant/0.025=0.014/0.025 =0.56

Dissolved Oxygen=(18*0.56*1000)/250

Dissolved Oxygen=40.32 mg/L